### v1

@@ 3.1 Minkowski spacetime
-- $(1-f(r_{s}))\frac{\sqrt{1-\alpha^{2}}}{r}+\sqrt{1-\alpha^{2}}\,y^{2}\frac{f^{\prime}(r_{s})}{r_{s}r}+\mathcal{O}(\frac{1}{r^{2}})$
$\rho=$ $-\frac{1}{16\pi}\left[\frac{1}{2}\left(v+\sqrt{1-\alpha^{2}}\right)^{2}\frac{f^{\prime}(r_{s})^{2}}{r_{s}^{2}}+\mathcal{O}(\frac{1}{r})\right]\,.$ (29)
Here, the first $\mathcal{O}(\frac{1}{r})$ term in $\theta$ is the background metric contribution and may be subtracted. Again the $(-ve)$ energy density is localized around the boundary of the warp bubble, even for the $\mathcal{O}(\frac{1}{r})$ terms which we omit writing down.
@@
$\left.+\mskip 1.5muv^{2}(x-x_{0})y^{2}(\alpha^{2}-1)\frac{f^{\prime}(r_{s})^{2}}{\alpha^{2}\mskip 1.5mur_{s}^{2}\mskip 1.5mur}+\mathcal{O}(r^{-2})\right]\,.$
@@ 3.2 AdS spacetime
where $H=\frac{1}{\ell}$ is the Hubble constant, $\ell$ is the AdS length scale and $\Lambda=-\frac{(D-1)(D-2)}{2\,\ell^{2}}$ is the cosmological constant. Following the arguments in section 2 one can construct the MP–charts with parameter $\epsilon$, however now the Eulerian observers are outgoing and travels with initial velocity $v_{\infty}$ from $r=0$ with proper time $T$ defined as
@@
Our focus is mainly on $\text{AdS}_{3}$ with the line element
@@
We shall now embed the warp drive centered at $x_{0}$ in this background. Closed form expressions for the $K_{ij}$ and $\rho$ can be obtained but as in the Minkowski case, we are mostly interested in the limit $x_{0}\gg(x-x_{0})\approx y\approx r_{s}\,$, where the warp drive is sufficiently far way from the singularity at the center $r=0$. As we have finite range charts, we have to treat the order of limits carefully. The radial distance $r$ in these MP–charts is cutoff at $r_{+}(\epsilon,H)\,$ therefore it is best to scale it as $r=(1-s_{r})\mskip 1.5mur_{+}$ such that $0\leq s_{r}\leq 1$ and then focus on the region near $s_{r}\approx 0\,$. To $\mathcal{O}(s_{r})$ and $\mathcal{O}(H)$, the expansion and energy density are as follows,
@@
Again, we notice the combination $\left(v-\sqrt{2\mskip 1.5mu(\epsilon^{2}-1)\mskip 1.5mus_{r}}\right)$ appearing at the zeroth order, however for AdS the background flow velocity and the warp drive velocity are both along $(+ve)$ $x$–direction. The first $\mathcal{O}(H)$ term in (37) is the contribution from the background AdS and maybe subtracted and the second term is very small. Similarly, in (38) the background contribution from the cosmological constant has been subtracted in the $\mathcal{O}(H^{2})$ terms. After the appropriate subtractions remaining terms are again proportional to $f^{\prime}(r_{s})^{\#}$ such that the change in $\theta$ and $\rho$ are localized around the boundary of the warp bubble. Interestingly, the zeroth order term in the $(-ve)$ energy density is suppressed by $(1/\epsilon^{2})$ thereby reducing the energy as $\epsilon>1$  [^11] This is also true for the Minkowski space. It is not visible in (3.1) as we truncated the expressions to $\mathcal{O}(v_{\infty})$..
@@
where $\alpha=1-4M$ and $0\leq\phi<2\pi$ DESER1984405 121212 $\text{dS}_{3}$ is bit tricky. In all three spacetimes, there is a second source at $r=\infty$. In Minkowski and $\text{AdS}_{3}$ spacetimes the geodesic distance between the source at $r=0$ and $r=\infty$ is infinite, so the second source can be ignored but for $\text{dS}_{3}$ it is finite and contributes.. The MP–charts with $\epsilon=\alpha$ doesn’t exist. Therefore, we focus on the $\epsilon=1,\,v_{\infty}=0$ charts given by
@@
We note that for physical particles, we have $0<\alpha\leq 1$. We shall now embed the warp drive in this conical $\text{AdS}_{3}$ background and do a similar analysis as done earlier. The important changes are that the spatial metric is now $\delta_{ij}$ and scaling of the radial coordinate is $r=(1-s_{r})\,{}^{c}r_{+}$ such that $0\leq s_{r}\leq 1$ and then focus on the region $s_{r}\approx 0$. The expansion $\theta$ is as given in (37) with the replacement $(\epsilon^{2}-1)\rightarrow(1-\alpha^{2})$. Parts of the $\mathcal{O}(H)$ not proportional to $f^{\prime}(r_{s})$ are from the background metric and should be subtracted. For the energy density $\rho$, the zeroth order $\mathcal{O}(H^{0})$ terms is the same as in (38) with the replacement $(\epsilon^{2}-1)\rightarrow(1-\alpha^{2})$ and omitting the $\epsilon^{2}$ in the denominator. From $\mathcal{O}(H)$, terms are different starting at $\mathcal{O}\left(\sqrt{s_{r}}\right)$. Again the background contribution from the cosmological constant needs to be subtracted from the $\mathcal{O}(H^{2})$ terms. Throughout, we also don’t have the $(1/\epsilon^{2})$ suppression as $\epsilon=1$.

### v2

@@ 3.1 Minkowski spacetime
$-(1-f(r_{s}))\frac{\sqrt{1-\alpha^{2}}}{r}+\sqrt{1-\alpha^{2}}\,y^{2}\frac{f^{\prime}(r_{s})}{r_{s}r}+\mathcal{O}\left(\frac{1}{r^{2}}\right)$
$\rho=$ $-\frac{1}{16\pi}\left[\frac{1}{2}\left(v+\sqrt{1-\alpha^{2}}\right)^{2}\frac{f^{\prime}(r_{s})^{2}}{r_{s}^{2}}+\mathcal{O}\left(\frac{1}{r}\right)\right]\,.$ (31)
Here, the first $\mathcal{O}\left(\frac{1}{r}\right)$ term in $\theta$ is the background metric contribution and may be subtracted. Again the $(-ve)$ energy density is localised around the boundary of the warp bubble, even for the $\mathcal{O}\left(\frac{1}{r}\right)$ terms which we omit writing down.
@@
$\left.+\mskip 1.5muv^{2}(x-x_{0})y^{2}(\alpha^{2}-1)\frac{f^{\prime}(r_{s})^{2}}{\alpha^{2}\mskip 1.5mur_{s}^{2}\mskip 1.5mur}+\mathcal{O}\left(\frac{1}{r^{2}}\right)\right]\,.$
@@
Energy considerations
So far, we have shown that the warp drives in the three dimensional Minkowski spacetime and its conical defect cousins in $\epsilon\geq 1$, MP–charts carry ($-ve$) energy densities localised around the warp bubble which means they violate the weak energy condition (WEC). Matt Visser et. al. in Santiago_2022 argued that for almost all Natário type warp drives discussed in the literature with a generic line element of the form,
$ds^{2}=-dt^{2}+\delta_{ij}\left(dx^{i}-v^{i}(\vec{x},t)dt\right)\left(dx^{j}-v^{j}(\vec{x},t)dt\right)\,,$ (34)
including the ones that claim ($+ve$) energy densities observed by the Eulerian observers violate either the strong energy condition (SEC), dominant energy condition (DEC), (WEC) or the null energy condition (NEC) for a generic observer. However, their analysis does not cover the warp drives discussed in this paper as the spatial metric $\gamma_{ij}\neq\delta_{ij}$. In general, the violation of the energy conditions follow the linkage Kontou_2020 ,
$\mathrm{NEC}\longrightarrow\mathrm{WEC}\longrightarrow\mathrm{DEC}\quad\text{and}\quad\mathrm{NEC}\longrightarrow\mathrm{SEC}\,.$ (35)
As we shall see, for the warp drives discussed here, (NEC) is violated which leads to the violation of the other energy conditions.
In order to explore the violation to the (NEC) it is best work with the natural co–moving orthonormal frame (triad) attached to the Eulerian observers of the the warp drive metric (14) with $N=1$. For the spatial metric $\gamma_{ij}$, we can choose the spatial dyad $\tilde{e}^{\hat{j}}_{i}$ such that $\gamma_{ij}=(\tilde{e}^{\hat{i}})_{i}(\tilde{e}^{\hat{j}})_{j}\delta_{\hat{i}\hat{j}}$ Then the triads for the metric $g_{\mu\nu}$ are
$(e^{\hat{0}})_{\mu}=u_{\mu}=(-1;0,0)\quad$ $\text{and}\quad(e^{\hat{i}})_{\mu}=(N^{j}(\tilde{e}^{\hat{i}})_{j};(\tilde{e}^{\hat{i}})_{i})$
$(e^{\hat{0}})^{\mu}=u^{\mu}=(1;-N^{i})\quad$ $\text{and}\quad(e^{\hat{i}})^{\mu}=(0;(\tilde{e}^{\hat{i}})^{i})\,.$ (36)
Usually it is assumed that the stress–energy tensor is of the Hawking–Ellis type I Hawking_Ellis_1973 which means it is diagonal in an orthonormal basis for the observer. Unfortunately, for the warp drive solutions we do not know a priori whether or not the stress–energy is Hawking–Ellis type I in general, nor do the Eulerian observers diagonalise the stress–energy tensor. This requires modification to the usual approach to testing energy conditions.
The null energy condition (NEC) demands that for all null vectors $k^{\mu}$, we must have $T_{\mu\nu}k^{\mu}k^{\nu}\geq 0$ In the orthonormal frame, let us take any two oppositely oriented null vectors $k_{+}^{\hat{\mu}}=(1,+k^{\hat{i}})$, and $k_{-}^{\hat{\mu}}=(1,-k^{\hat{i}})$ where $k^{\hat{i}}$ is an arbitrary unit spatial 2–vector. Then the (NEC) would demand,
$T_{\hat{\mu}\hat{\nu}}k_{+}^{\hat{\mu}}k_{+}^{\hat{\nu}}$ $=\rho+2T_{\hat{0}\hat{i}}k^{\hat{i}}+T_{\hat{i}\hat{j}}k^{\hat{i}}k^{\hat{j}}\geq 0$ (37)
$T_{\hat{\mu}\hat{\nu}}k_{-}^{\hat{\mu}}k_{-}^{\hat{\nu}}$ $=\rho-2T_{\hat{0}\hat{i}}k^{\hat{i}}+T_{\hat{i}\hat{j}}k^{\hat{i}}k^{\hat{j}}\geq 0\,,$
which upon averaging implies the following one–way relation, $\mathrm{(NEC)}\longrightarrow\rho+T_{\hat{i}\hat{j}}k^{\hat{i}}k^{\hat{j}}\geq 0$ For a dyad $k_{A}^{\hat{i}}$ of two mutually orthogonal unit vectors, for each $A$ (NEC) implies, $\rho+T_{\hat{i}\hat{j}}k_{A}^{\hat{i}}k_{A}^{\hat{j}}\geq 0$ Noting that by construction $\sum_{A}k_{A}^{\hat{i}}k_{A}^{\hat{j}}=\delta^{\hat{i}\hat{j}}$ and averaging over the two members of the dyad gives
$\rho+T_{\hat{i}\hat{j}}\left(\frac{1}{2}\sum_{A}k_{A}^{\hat{i}}k_{A}^{\hat{j}}\right)\geq 0\quad\text{or}\quad\rho+\frac{1}{2}T_{\hat{i}\hat{j}}\delta^{\hat{i}\hat{j}}\geq 0\,.$ (38)
Even if the 2–stress tensor $T_{\hat{i}\hat{j}}$ is not diagonal, we can define the average pressure as usual, $\bar{p}=\frac{1}{2}T_{\hat{i}\hat{j}}\delta^{\hat{i}\hat{j}}$ Finally, we have the one–way relation,
$\mathrm{(\textbf{NEC})}\longrightarrow\rho+\bar{p}\geq 0\,.$ (39)
Similarly, for the other energy conditions, the implications are  [^10] There are other energy conditions like trace energy condition (TEC), flux energy condition (FEC) etc. and various weaker average energy conditions of which averaged null energy condition (ANEC) is widely applicable.
  • $\bullet$
    (WEC) : $\rho\geq 0$ and $\rho+\bar{p}\geq 0$ coming from the demand $T_{\mu\nu}\ell^{\mu}\ell^{\nu}\geq 0$ for all timelike vectors $\ell^{\mu}$.
  • $\bullet$
    (SEC) : $\rho+2\bar{p}\geq 0$ and $\rho+\bar{p}\geq 0$ coming from the demand $\left(T_{\mu\nu}-\frac{1}{2}Tg_{\mu\nu}\right)\ell^{\mu}\ell^{\nu}\geq 0$ for all timelike vectors $\ell^{\mu}$.
  • $\bullet$
    (DEC) : $|\bar{p}|\leq\rho$ coming from the demand $T_{\mu\nu}\ell^{\mu}\tilde{\ell}^{\nu}\geq 0$ for all future–pointing timelike vectors $\ell^{\mu}$ and ~$\tilde{\ell}^{\nu}$.
To compute the average pressure $\bar{p}=\frac{1}{2}T_{\hat{i}\hat{j}}\delta^{\hat{i}\hat{j}}=\frac{1}{2}T_{ij}\gamma^{ij}=\frac{1}{2}T_{\mu\nu}P^{\mu\nu}$ for the warp drives discussed in this paper we can use (23) along with other Gauss–Codazzi relations,
$R_{ij}$ $={}^{(2)}R_{ij}+\mathcal{L}_{u}K_{ij}-2(K^{2})_{ij}+KK_{ij}$ (40)
$R_{uu}$ $=-\mathcal{L}_{u}K-\mathrm{tr}(K^{2})$
$R_{ui}$ $=D^{k}K_{ik}-D_{i}K\,,$
to evaluate the Einstein tensor $G_{ij}$ and hence $T_{ij}$ For the warp drive described by (14) and (15) to (18), the energy density is already evaluated in (27). To check for the violation of (NEC) to the leading order in $r$, we note that the form function $f(r_{s})$ is a monotonically decreasing function such that $f^{\prime}(r_{s})<0$, $f^{\prime\prime}(r_{s})<0$ and $0\leq f(r_{s})\leq 1$ Therefore, (NEC) is violated for the $\epsilon\geq 1$, MP–charts warp drives,
$\rho+\bar{p}=$ $-\frac{1}{32\pi\epsilon^{2}r_{s}^{3}}\Bigl{[}\left(v+\sqrt{\epsilon^{2}-1}\right)^{2}\left\{y^{2}\epsilon^{2}(f(r_{s})-1)f^{\prime}(r_{s})\right.$
$\left.+r_{s}\left(y^{2}+\epsilon^{2}(x-x_{0})^{2}\right)f^{\prime}(r_{s})^{2}\right.$ (41)
$\left.+r_{s}\epsilon^{2}(x-x_{0})^{2}(f(r_{s})-1)f^{\prime\prime}(r_{s})\right\}\Bigr{]}\leq 0\,.$
We can do similar exercises for the warp drives embedded in the conical defect backgrounds, again (NEC) is violated as expected. The (NEC) violation for these warp drive solutions leads to violations of other energy conditions, (WEC), (DEC) and (SEC).
We now turn our focus to give a reasonable estimate of the total energy measured by the Eulerian observers. For the $\epsilon\geq 1$, MP–charts warp drives, shifting the origin to the location of the warp drive and to the leading order in $r$, the total energy is given by integrating the energy density (27) over a spatial slice,
$E$ $=\int T_{\mu\nu}u^{\mu}u^{\nu}d^{2}x$ (42)
[…]
Of course, $E$ is ($-ve$) and depends on the choice of the form function $f(r)$. Instead of choosing the usual $f(r)$ (19), we shall use the piecewise–continuous form function suggested by Pfenning and Ford Pfenning_1997 as it provides better estimates for the total energy,
$f(r)=\begin{cases}1&r\leq R-\frac{\delta}{2}\\ \frac{1}{2}+\frac{R-r}{\delta}&r\in\left(R-\frac{\delta}{2},R+\frac{\delta}{2}\right)\\ 0&r\geq R+\frac{\delta}{2}\end{cases}$ (43)
Here, $R$ is the radius of the warp drive bubble and $\delta$ is its wall thickness. Plugging this into (42), we get
[…]
$=-\frac{\left(v+\sqrt{\epsilon^{2}-1}\right)^{2}}{64\pi\epsilon^{2}}\frac{R}{\delta}\,.$
We can choose a reasonable size for the bubble, say $R=100\mathrm{~{}m}$ and make a choice for the background flow velocity ($v_{\infty}$) but the crucial ingredient is to estimate the maximum thickness of the bubble wall so as to reduce the overall ($-ve$) energy to its minimum.
The restrictions to the thickness of the bubble can be obtained from quantum inequalities (QI) which do allow ($-ve$) energies but places serious limitations on its magnitude and duration PhysRevD.43.3972 . It would be very difficult to quantise a scalar field on a warp drive background and arrive at the exact (QI), but upto corrections in the inverse powers of the local radius of curvature, the flat spacetime (QI) can be applied to curved spacetimes with the additional restriction that the ($-ve$) energy be sampled on timescales smaller than the minimum local radius of curvature Pfenning_1997_1 . Most of the literature on (QI) is in four dimensions but the warp drives in this paper are in three dimensions. Redoing the computations, we arrive at the expression,
[…]
where $g(T)=\frac{T_{0}}{\pi(T^{2}+T_{0}^{2})}$ is a suitable sampling function. For the $\epsilon\geq 1$, MP–charts warp drives in consideration we notice that the largest component of the Riemann tensor is given by
$\bigl{|}R_{\hat{T}\hat{y}\hat{T}\hat{y}}\bigr{|}=\frac{3\left(v+\sqrt{\epsilon^{2}-1}\right)^{2}y^{2}}{4\epsilon^{2}r^{2}}f^{\prime}(r)^{2}\,,$ (46)
which yields the minimum radius of curvature to be
$r_{\min}\equiv\frac{1}{\sqrt{\bigl{|}R_{\hat{T}\hat{y}\hat{T}\hat{y}}\bigr{|}}}\sim\frac{2\epsilon\delta}{\sqrt{3}\left(v+\sqrt{\epsilon^{2}-1}\right)}\,,$ (47)
near the boundary of the bubble. As mentioned earlier, the sampling time must be smaller than this length scale, so we can take
$T_{0}=\Delta\frac{2\epsilon\delta}{\sqrt{3}\left(v+\sqrt{\epsilon^{2}-1}\right)}\quad 0<\Delta\ll 1\,,$ (48)
where $\Delta$ is an unspecified parameter which we choose to be, $\Delta=1/10$ Following the steps in Pfenning_1997 , (QI) puts a bound on the thickness of the bubble,
$\delta\leq\frac{3\sqrt{3}}{4\epsilon\Delta^{3}}\left(v+\sqrt{\epsilon^{2}-1}\right)\sim 10^{3}\frac{\left(v+\sqrt{\epsilon^{2}-1}\right)}{\epsilon}\ell_{p}\,,$ (49)
where $\ell_{p}$ is the Planck length  [^11] In three dimensions, $\ell_{p}=\frac{\hbar G_{3}}{c^{3}}$ but if we assume that we reach three dimensions by compactifying one spatial direction of order $\ell_{p}$ in four dimensions, then $G_{3}=G_{4}/\ell_{p}$ and $\ell_{p}=1.6\times 10^{-35}\mathrm{~{}m}$.. We can assume the mass of a Milky–way galaxy in three dimensional universe to be two third of the actual Milky–way, $M_{\mathrm{Milky}}=1.3\times 10^{33}~{}m_{p}$ Putting it all together in (44), we get
$E\leq-3\times M_{\mathrm{Milky}}\frac{\left(v+\sqrt{\epsilon^{2}-1}\right)}{\epsilon}\,.$ (50)
It is about $10^{20}$ times better than the four dimensional result but still a stupendously large amount of ($-ve$) energy is required to operate the warp drive in three dimensions.
The energy considerations for the warp drives embedded in the conical defect backgrounds are similar to the above analysis. If we compare with the original $\epsilon=1$, Alcubierre warp drive in three dimensions, the energy density and total energy of the $\epsilon=1$, conical defect warp drives scales by a factor of $\left(\frac{\left(v+\sqrt{1-\alpha^{2}}\right)}{v}\right)\geq 1$, assuming that the warp drive is travelling outwards. This is expected as the conical defect background has a flow inwards towards the origin. Similarly, for the other case of $\epsilon=\alpha$, the suppressing ratio is $\frac{1}{\alpha^{2}}\leq 1$ i.e. it has the potential to lower the energy requirements, but unfortunately it also requires ($-ve$) mass to create the defect itself. The analysis also goes through for the warp drives embedded in the AdS/dS and their conical defect cousins discussed in the next section in the small Hubble constant, $H\ll 1$ limit  [^12] The finite $H$ computations are challenging. Firstly the metric is complicated in the Cartesian coordinates which is necessary to correctly embed the warp drives and secondly the (QI) for AdS/dS spaces would require quantising scalar fields in curved spacetimes. We shall attempt a more complete analysis for these cases in a future work..
@@ 3.2 AdS spacetime
where $H=\frac{1}{\ell}$ is the Hubble constant, $\ell$ is the AdS length scale and $\Lambda=-\frac{(D-1)(D-2)}{2\,\ell^{2}}$ is the cosmological constant. Following the arguments in section 2 one can construct the MP–charts with parameter $\epsilon$, however now the Eulerian observers are outgoing and travels with initial velocity $v_{\infty}$ from $r=0$ with proper time $T$ defined as,
@@
Our focus is mainly on $\text{AdS}_{3}$ with the line element,
@@
We shall now embed the warp drive centred at $x_{0}$ in this background. Closed form expressions for the $K_{ij}$ and $\rho$ can be obtained but as in the Minkowski case, we are mostly interested in the limit $x_{0}\gg(x-x_{0})\approx y\approx r_{s}\,$, where the warp drive is sufficiently far way from the singularity at the centre $r=0$. As we have finite range charts, we have to treat the order of limits carefully. The radial distance $r$ in these MP–charts is cutoff at $r_{+}(\epsilon,H)\,$ therefore it is best to scale it as $r=(1-s_{r})\mskip 1.5mur_{+}$ such that $0\leq s_{r}\leq 1$ and then focus on the region near $s_{r}\approx 0\,$. To $\mathcal{O}(s_{r})$ and $\mathcal{O}(H)$, the expansion and energy density are as follows,
@@
We notice the combination, $\left(v-\sqrt{2\mskip 1.5mu(\epsilon^{2}-1)\mskip 1.5mus_{r}}\right)$ appearing at the zeroth order, however for AdS the background flow velocity and the warp drive velocity are both along $(+ve)$ $x$–direction. The first $\mathcal{O}(H)$ term in (56) is the contribution from the background AdS and maybe subtracted and the second term is very small. Similarly, in (57) the background contribution from the cosmological constant has been subtracted in the $\mathcal{O}(H^{2})$ terms. After the appropriate subtractions remaining terms are again proportional to $f^{\prime}(r_{s})^{\#}$ such that the change in $\theta$ and $\rho$ are localised around the boundary of the warp bubble. Interestingly, the zeroth order term in the $(-ve)$ energy density is suppressed by $(1/\epsilon^{2})$ thereby reducing the energy as $\epsilon>1$  [^14] This is also true for the Minkowski space. It is not visible in (27) as we truncated the expressions to $\mathcal{O}(v_{\infty})$..
@@
where $\alpha=1-4M$ and $0\leq\phi<2\pi$ DESER1984405 151515 $\text{dS}_{3}$ is bit tricky. In all three spacetimes, there is a second source at $r=\infty$. In Minkowski and $\text{AdS}_{3}$ spacetimes the geodesic distance between the source at $r=0$ and $r=\infty$ is infinite, so the second source can be ignored but for $\text{dS}_{3}$ it is finite and contributes.. The MP–charts with $\epsilon=\alpha$ doesn’t exist. Therefore, we focus on the $\epsilon=1,\,v_{\infty}=0$ charts given by
@@
We note that for physical particles, we have $0<\alpha\leq 1$.
We shall now embed the warp drive in this conical $\text{AdS}_{3}$ background and do a similar analysis as done earlier. The important changes are that the spatial metric is now $\delta_{ij}$ and scaling of the radial coordinate is $r=(1-s_{r})\,{}^{c}r_{+}$ such that $0\leq s_{r}\leq 1$ and then we focus on the region $s_{r}\approx 0$. The expansion $\theta$ is as given in (56) with the replacement $(\epsilon^{2}-1)\rightarrow(1-\alpha^{2})$. Parts of the $\mathcal{O}(H)$ not proportional to $f^{\prime}(r_{s})$ are from the background metric and should be subtracted. For the energy density $\rho$, the zeroth order $\mathcal{O}(H^{0})$ terms is the same as in (57) with the replacement $(\epsilon^{2}-1)\rightarrow(1-\alpha^{2})$ while omitting the $\epsilon^{2}$ in the denominator. From $\mathcal{O}(H)$, terms are different, in this case starting at $\mathcal{O}\left(\sqrt{s_{r}}\right)$. Again the background contribution from the cosmological constant needs to be subtracted from the $\mathcal{O}(H^{2})$ terms. Throughout, we also don’t have the $(1/\epsilon^{2})$ suppression as $\epsilon=1$.